A box contains 10 calculators, 5 of which are defective. If 4 are selected at random,find the following probabilities:
A.exactly 2 are defective
B.at least 1 is defective

Answer by stanbon(75887)

A box contains 10 calculators, 5 of which are defective. If 4 are selected at random,find the following probabilities:
----------------------------
A. P(exactly 2 are defective)
# of ways to succeed:: 5C2*5C2 = 10*10 = 100
# of possible groups of 4 = 10C4 = 210
P(exactly 2 defective) = 100/210 = 10/21
--------------------------------------------------------
B.at least 1 is defective = 1 - P(none are defective)
= 1 - [5C4/10C4] = 1 - (5/210) = 205/210 = 41/42
-------------------
Cheers,
Stan H.